# Quantitative analysis of the value of investment in research facilities, with examples from cyberinfrastructure

### v2

@@ Abstract
Results: We present methods developed primarily for analyses of cyberinfrastructure. Most analyses comparing investment in university-based cyberinfrastructure facilities with purchasing services from commercial sources demonstrate positive results for economic and scientific research. A recent assessment, based on a comprehensive accounting approach, has shown that for one large publicly funded cyberinfrastructure project the value delivered to the USA economy and society exceeded the cost to USA taxpayers [1].
@@ Introduction
Institutions of higher education are called on to educate, to engage in research, to create (scientific breakthroughs, works of art), and to contribute to the public good. Every college and university, whether public or private, has its own mission and priorities attached to these core activities. Since the beginning of the modern era of research activities, which in the USA began with the 1945 report by Vannevar Bush entitled “Science, the Endless Frontier” [2], there have been very few times during which some universities were able to spend money at will. During most of the past 80 years, availability and allocation of funding has been a matter of some concern to almost all institutions of higher education. This paper presents longstanding and new methods for the evaluation of investment in research-enabling facilities, based particularly on methods developed for analysis of computing infrastructure used to enable research activities. All of these methods are in principle extendable to investment in research generally.
@@
In 2025 a new sort of financial challenge has emerged. The USA federal government has threatened dozens of institutions of higher education with the loss of grant funds due to alleged violations of the USA Civil Rights Act of 1964. Some universities are at risk of having hundreds of millions to billions of dollars in total grant funding canceled [6]. Nothing in this paper will be relevant to the resolution of actual legal misdeeds. However, if, as some have asserted, the lodging of these allegations rests on an underlying ideological questioning of the value of higher education in the USA, then demonstration of benefits to the USA as a whole that derive from research conducted at institutions of higher education may be of use in discussions. This is not just a problem in the USA. Although this particular legal issue may be specific to the USA, ideological questions about the value of higher education are not.
@@
National policies related to research at institutions of higher education are tied to success in the commercialization of new technology products that improve quality of life, global competitiveness and defense, and profits [14] [15]. Leaders at institutions of higher education are at the vanguard of a technology ecosystem that may take decades to turn an idea into a widely used product. Because it may take decades for a new discovery to translate into a product or service that improves lives, stakeholders of all sorts need good information through this many-year innovation cycle so that projects may be fairly evaluated on an ongoing basis between the points of “initial investments” and “useful end product that demonstrates the value of investments.” It is also important to note that while the methods presented here arise from the current challenges being faced in higher education in highly developed countries, they are equally relevant in justifying expansion of research activities in academia in the global south.
@@
In the remainder of this paper, we summarize recent developments in the quantitative assessment of the value of research facilities, with a particular emphasis on methods recently developed in studies of cyberinfrastructure. We briefly consider some of the implications of the recent results of particular studies of the value of cyberinfrastructure. Although we do not present any new data, reconsidering the result of five recent publications of the authors together yields new insights. We then discuss the extension of the methods to other kinds of research-enabling facilities and research initiatives generally. The methods presented here were developed for evaluation of cyberinfrastructure in a higher education setting, but are applicable to research and development in any setting. We show an example of how the now decades-old Balanced Scorecard concept [16] may be adapted to concisely convey quantitative metrics relevant to research-enabling facilities and research generally. We then conclude with a few comments about potential areas of future research.
@@
Different kinds of readers are likely to draw different insights from this paper. We hope that leaders and staff of research-enabling facilities and researchers in general will see new ways to express the value of their activities. We hope that a skeptical CFO (Chief Financial Officer) might take this paper to a colleague in economics or accounting and receive back the assessment, “within the bounds of the data that are available, the methodology and conclusions are reasonable.” We hope that any senior leader in higher education will find, for purposes of communicating outside their home institution, new means to explain the value of their institutions to important stakeholders and, for purposes within their home institution, helpful approaches for financial and strategic decision making. We dearly hope that researchers and higher education administrators in the global South will find tools that are applicable in their contexts to aid their work. Last and not least we hope that researchers in the areas of economics of innovation and science of science will find here sources for fresh ideas in these important areas.
@@
Economic studies that attempt to draw a general link between research and economic benefit go back more than a century, to the 1912 publication of Schumpeter’s Theorie der wirtschaftlichen Entwicklung (Theory of Economic Development) [17]. Qualitative analyses have long pointed out relationships between investment in research and economic growth. A current example of this sort of approach, used in the USA since at least the 1990s, can be found in [18]. This sort of analysis links broad areas of economic activity, such as telecommunications, back to specific suites of USA federal government investment in research. While these qualitative relationships are convincing, they are drawn out only at very large scales and only in rare cases point back to a single research-enabling facility as a causal factor. As a result, such approaches are not generally helpful in evaluating a particular research-enabling facility and not at all helpful in managing budgets within institutions of higher education.
@@
In 2016, Del Bo reviewed the rate of return of large-scale “research infrastructures,” which is a term from the literature in the European Union [20]. The term ”research infrastructure” is derived from Florio’s early work and is defined as “high-capital intensity and long-lasting facilities and equipment, typically operating in oligopoly conditions, whose objective is to support economic development and produce social benefits through the generation of new knowledge and, often, other spillover effects” [21]. Del Bo’s analysis of return on investment (ROI) in research infrastructures helpfully creates a categorization of studies based on scope, ranging from macroeconomic to project-level studies [20]. Del Bo’s excellent review of macroeconomic methods applicable to research infrastructures is based largely on endogenous growth models. Del Bo’s efforts were focused on the European Union, but the methods outlined are broadly applicable to any situation involving an appropriately large and long-lasting research infrastructure.
The European facilities of CERN (Conseil européen pour la recherche nucléaire - the Europe-based organization for Nuclear Research) - is perhaps the best known example of a persistant and large-scal Research Infrastructure in the sense of Del Bo’s definition. The investments in CERN are so large that it is possible to assess their impacts through national and multinational macroeconomic studies. Florio followed his early paper-length analyses with an extensive book devoted to “social cost-benefit of research infrastructures” [22]. This book offers a thorough conceptual overview of types of benefit that may derive from investment in large-scale Research Infrastructures and quantitative methods for estimating such effects. The summary of all of the research of the public value of public investment in CERN is that it has a net positive impact on society. Many other of the Research Infrastructures reviewed by Del Bo also have a net positive economic impact.
@@
CERN and cyberinfrastructure facilities at universities in the USA have been the subject of more cost-benefit and ROI analyses than most - perhaps any - research-enabling facilities in the world. Most cyberinfrastructure facilities involve orders of magnitude less investment and less economic impact. As a result their impact is generally not detectable via the macroeconomic approaches used in analysis of CERN and other large-scale Research Infrastructures. Recognizing the extensive use of the term Research Infrastructure as defined by Florio and implying persistent and very large-scale facilities, we will focus on activities at a smaller scale. In particular we will use the term research-enabling facility to mean just what those words appear to mean limited to facilities at scales smaller than a Research Infrastructure as defined by Florio.
In this section, we present project-level approaches to quantitative assessment of investment in cyberinfrastructure. Much of the research cited in this current paper is project-level analysis of one of the largest cyberinfrastructure projects ever funded by the USA’s National Science Foundation (NSF), called XSEDE - the eXtreme Science and Engineering Discovery Environment. XSEDE included supercomputers, visualization resources, data storage facilities, and training and support services, described in detail in [26] and [11]. XSEDE operated from 2011 to 2022, providing services supporting non-classified research by the USA national research community. Because of its scale and visibility, the value provided to the USA by XSEDE has been examined with particular care.
Our review of project-level approaches to evaluating research-enabling facilities uses a categorization based on that presented by Del Bo and expanded to include other and newer analytic methods:
@@
We focus particularly on the critical information that potentially skeptical stakeholders, including government officials and the general public, commonly ask for: Is the work enabled by a research-enabling facility sufficiently financially significant that the investment in that facility is financially worthwhile and beneficial to society as a whole?
@@
The total quality management movement reached the USA with force in the 1980s. Total quality management is “a term first used to describe a management approach to quality improvement” [27] and is explained in detail in [28]. One focus of total quality management is the creation and measurement of a set of metrics that meaningfully define quality and costs for an organization. Such metrics have long included client- or user- reported satisfaction, total cost of ownership (TCO), and usage metrics that are typically specific to organizational type and an organization’s specific goals. Such approaches have been applied to computing services in academia since at least the 1990s [29]. All of these metrics are useful but have limitations.
Client satisfaction surveys are a basic aspect of the quality assessment of any research-enabling infrastructure or facility. (Term usage varies. “Client” is regarded as a respectful term among many who are steeped in TQM methodologies, although the term “user” is also commonly used.) Highly negative survey ratings of course indicate problems. Positive ratings demonstrate that the users of a facility like it, which is good. However, positive satisfaction scores do not necessarily prove that a facility is a cost-effective or a strategic investment [29].
@@
For many categories of research-enabling facility there are also specific sorts of traditional operational metrics. There may also be particular goals for a specific organization tied to its strategic objectives. Typical metrics for the operation of cyberinfrastructure facilities include percentage uptime (time during which a system was available for use as a percentage of theoretical maximum), the number of jobs submitted and completed, CPU- and GPU- hours used per unit, and the amount of data stored. Again, all of these metrics must be carefully interpreted. Unusually low usage numbers relative to resource availability may be indicative of some sort of problem - perhaps there has been insufficient training, or the system is not perceived by users to match their needs. Low usage statistics may also be found immediately after a large upgrade to the capability of computational systems when the people who use the system have not adjusted to the increased availability of cycles. Consistent usage near peak capacity may be indicative of a balanced match of availability and need. Consistent usage at peak capacity may indicate inadequate supply, or a doctoral student iteratively reanalyzing their data to avoid actually writing their dissertation, or a faculty member running mindless parameter sweeps rather than running carefully thought out and efficient computational experiments. (These are not hypothetical examples; the last named author was responsible for HPC systems for 30 years and experienced every one of these examples at least once.)
There will be discipline-specific operational metrics appropriate to any sort of research-enabling facility. While such metrics can clarify the activities and utilization of a research-enabling facility among practitioners in a given area,these operational metrics are not easily translatable into terms that relate value to investment. As such they are unlikely to resonate intuitively with certain important categories of stakeholders, potentially ranging from taxpaying voters to elected officials to University Presidents and CFOs (Chief Financial Officers).
@@
One example of the analysis of particular sorts of research facilities is presented in von Laszewski et al. [32], who performed a bibliometric analysis of peer-reviewed scientific publications created with analyses of simulations performed on XSEDE cyberinfrastructure facilities. They investigated the question: Are research publications created through the use of XSEDE cited more often than similar papers created without the use of XSEDE? Their analysis suggested that the answer was yes. Unfortunately, however, this analysis suffered from a significant conceptual shortcoming. To investigate the impact on research of one particular type of research-enabling facility, one really ought to ask two questions: 1) Are papers produced using a certain type of research-enabling facility more impactful, in general, than similar papers produced without this type of research-enabling facility?; and 2) Within the category of papers produced using a particular category of research-enabling facility, are those produced using one particular facility (or perhaps one type of instrument) more impactful than other papers produced using facilities of the same type (or instruments of the same category)? The line of analysis suggested by von Laszewski et al. is interesting, but a more careful analysis is needed, hopefully published in one of the standard journals that focus on bibliometrics or the science of science.
@@
Empirically derived econometric estimators can be used to estimate the direct and indirect economic effects of investments in many kinds of projects and services. The Regional Input–Output Modeling System (RIMSII) uses empirically derived multipliers to estimate the total direct and indirect economic impact and total number of jobs created by many sorts of investments [34]. These multipliers are calculated for a variety of kinds of investment, ranging from building a parking garage to investing in jobs and facilities at a university, for specific geographic regions. IMPLAN (originally short for “impact analysis for planning”) offers the same approach globally [35]. Because these approaches use general categories of investment, they may underestimate the value of investment in highly unusual areas. For example, when estimating the value of investment in advanced cyberinfrastructure systems within the USA, one could potentially use a RIMSII multiplier from one of two types of “Professional, scientific, and technical services,” or one of two types of “Junior colleges, colleges, universities, and professional schools.” The impact of some specialized research-enabling facility might well be greater than the calculated average for such relatively broad categories. However, these estimation methods are well-known and widely trusted. They produce figures that constitute an unimpeachable lower bound for the overall economic impact of any investment.
@@
One of the earliest concepts in accounting for the value of a particular instance of publicly funded research-enabling facility has been referred to as “leverage.” For publicly funded research facilities, leverage is simply the financial value of the sum of grant awards received by the publicly funded research projects that make use of that facility. This total value can be useful in concise statements explaining the value of a facility. Using data from [1] about how much the USA federal government invested in XSEDE and how much it invested in other grant awards that used XSEDE resources, for example, one could say that, “The NSF leveraged USD 200M of investment in XSEDE to support USD 4.5B in total grant awards that made use of XSEDE facilities and services.” These figures must be used with a measure of care, however, because of the matter of shared credit. Suppose that a discovery critically depends on two different research-enabling facilities. The value of such a discovery can be considered as “leverage” for each of those facilities. But it would be double counting to count the full value as part of the “return on the investment” in each facility, something that has been done in some reports. It is certainly unlikely in general to be wholly true that research conducted with the use of a particular facility could or would not have been successfully carried out without the use of that facility. Such a claim might be true in the case of the research conducted with the Large Hadron Collider that led to the verification of the existence of the Higgs Boson. But in general one is well advised to be wary of contrapositive statements of the form “without research-enabling facility Z, discovery Beta would not have been possible.” Certainly in the case of many kinds of research-enabling facilities, some of the research done using that facility might well have been successfully carried out by clever scientists using some other facility. Some other research might have occurred, but not as efficiently or as well. And yet other research may indeed not have happened at all. So, while a “leverage” total can be useful shorthand, it is necessary to be careful when using this term.
@@
This formula is essentially a way to calculate the financial efficiency of providing a service. In terms of public sector research and university activities, this formula shows whether something is being done at greater or lesser cost than would be experienced if products or services were simply bought on the open market. ROIproxy is a way to measure the financial effectiveness resulting from investments in cyberinfrastructure over any particular time period, in the absence of anything being bought or sold. In this formula, as in the formula for ROI, a value of 1.0 indicates a break-even level. A value greater than 1.0 indicates that the value of services provided when calculated based on market costs exceeds the actual cost of operating the facility. These calculations are conceptually straightforward. The denominator is the TCO for the relevant period of time. For the numerator, one tallies as many services and products as possible that were created using the cyberinfrastructure resources, finds the prevailing market prices for as many of these services as possible, and arrives at a total estimated market value for the relevant time period. Credible valuations are available for a surprising array of entities, including researcher time, scientific accomplishments such as a peer-reviewed publication or Nobel Prize, educational outcomes, and even improvements in quality of life. These valuations may be done annually or over longer time periods. The results are approximate, of course. Calculated results are likely to be an underestimation of the actual long-term value. There will often be benefits that are impossible to quantify and thus not included in the calculations. Further, the economic value of some benefits will increase over time as the scope of application expands. Examples include health care and other benefits to people’s quality of life.
@@
A value >1 favors on-premises as the more economical choice.
@@
The International Integrated Reporting Framework is of use in answering this more difficult and far more important question. The purpose of the International $<$IR$>$ Framework (as it is typically referred to) is “to explain to providers of financial capital how an organization creates, preserves, or erodes value over time” [37]. This framework includes six categories, intended to capture all possible types of value creation: financial capital, manufactured capital, intellectual capital, human capital, social and relationship capital, and natural capital. This structure, according to the IFRS (International Financial Reporting Standards) website, benefits all stakeholders, including “legislators, regulators and policy-makers.” The IFRS has a grand goal of transforming corporate reporting, so the description of their reporting framework extends well beyond what is needed for an assessment of the value of an instance of a research-enabling facility. However, the comprehensiveness of the framework, its clarity, and its backing and promotion by an international consortium of accountancy researchers and experts provide great credibility to the portion relevant to the evaluation of research facilities and research projects generally (pp. 18-23 in [37]).
@@
Within academia, it can be useful to quantitatively examine important outputs even if it is not immediately possible to associate a financial value with them. And there are indeed non-financial metrics within academia that are important in an institution’s reputation, ranking, and its recruitment of students, faculty, and staff. In a 1928 work, Cobb and Douglas proposed a “production function” that expressed factory output (Y) as a function of investment in capital (K) and labor (L) [38]. Smith built on this approach to relate the production of a variety of academic outputs (Y) to capital investment in systems (as measured in TeraFLOPS) and labor costs for the staff that operate and support them as[39]:
@@
Table 1: Relationship of academic output metrics to on-premise HPC TeraFLOPS capacity and investment in staff supporting use of HPC systems at Purdue University, 1999–2020.
@@
Table 2: Percent variance in academic output metrics at Purdue University, explained by on-premise HPC TeraFLOPS capacity and investment in staff supporting use of HPC systems at Purdue University, 1999–2020.
@@
Table 3: Relationship of academic output metrics to on-premise HPC TeraFLOPS capacity and investment in staff supporting use of HPC systems at 5 R1 universities.
@@
Table 4: Percent variance in academic output metrics explained by on-premise HPC TeraFLOPS capacity and investment in staff supporting use of HPC systems at 5 R1 universities.

### v3

@@ Abstract
Results: We present methods developed primarily for analyses of cyberinfrastructure. Most analyses comparing investment in university-based cyberinfrastructure facilities with purchasing services from commercial sources demonstrate positive results for economic and scientific research. For example, one recent analysis showed a positive relationship between investment in cyberinfrastructure facilities and high-impact publications and number of PhDs conferred by a university. Another assessment showed that a publicly funded cyberinfrastructure project delivered to the USA economy and society exceeded the cost to USA taxpayers.
@@ Introduction
Institutions of higher education are called on to educate, to engage in research, to create (scientific breakthroughs, works of art), and to contribute to the public good. Every college and university, whether public or private, has its own mission and priorities attached to these core activities. Since the beginning of the modern era of research activities, which in the USA began with the 1945 report by Vannevar Bush entitled “Science, the Endless Frontier” [1], there have been very few times during which some universities were able to spend money at will. During most of the past 80 years, availability and allocation of funding has been a matter of some concern to almost all institutions of higher education. This paper presents longstanding and new methods for the evaluation of investment in research-enabling facilities, specifically focusing on methods developed for analysis of computing infrastructure used to enable research activities. All of these methods are in principle extendable to investment in research generally.
@@
In 2025 a new sort of financial challenge has emerged. The USA federal government has threatened dozens of institutions of higher education with the loss of grant funds due to alleged violations of the USA Civil Rights Act of 1964. Some universities are at risk of having hundreds of millions to billions of dollars in total grant funding canceled [5]. Nothing in this paper will be relevant to the resolution of actual legal misdeeds. However, if, as some have asserted, the lodging of these allegations rests on an underlying ideological questioning of the value of higher education in the USA, then demonstration of benefits to the USA as a whole that derive from research conducted at institutions of higher education may be of use in discussions. This is not just a problem in the USA. Although this distinct legal issue may be specific to the USA, ideological questions about the value of higher education are not.
@@
National policies related to research at institutions of higher education are tied to success in the commercialization of new technology products that improve quality of life, global competitiveness and defense, and profits [14] [15]. Leaders at institutions of higher education are at the vanguard of a technology ecosystem that may take decades to turn an idea into a widely used product. Because of the time lag between a new discovery and its translation into improving lives, stakeholders of all sorts need good information to enable ongoing evaluation during this development and translation process. It is also important to note that while the methods presented here arise from the current challenges being faced in higher education in highly developed countries, they are equally relevant in justifying expansion of research activities in academia in the Global South.
@@
In the remainder of this paper, we summarize recent developments in the quantitative assessment of the value of research facilities, with an emphasis on methods recently developed in studies of cyberinfrastructure. We briefly consider implications for the recent results of studies focused on the value of cyberinfrastructure. Although we do not present any new data, reconsidering the result of five recent publications of the authors together yields new insights. We then discuss the extension of the methods to other kinds of research-enabling facilities and research initiatives generally. The methods presented here were developed for evaluation of cyberinfrastructure in a higher education setting, but are applicable to research and development in any setting. We show an example of how the now decades-old Balanced Scorecard concept [16] may be adapted to concisely convey quantitative metrics relevant to research-enabling facilities and research generally. We then conclude with a few comments about potential areas of future research.
@@
Different kinds of readers are likely to draw different insights from this paper. We hope that leaders and staff of research-enabling facilities and researchers in general will see new ways to express the value of their activities. We hope that a skeptical CFO (Chief Financial Officer) might take this paper to a colleague in economics or accounting and receive back the assessment, “within the bounds of the data that are available, the methodology and conclusions are reasonable.” We hope that any senior leader in higher education will find, for purposes of communicating outside their home institution, new means to explain the value of their institutions to important stakeholders and, for purposes within their home institution, helpful approaches for financial and strategic decision making. We dearly hope that researchers and higher education administrators in the Global South will find tools that are applicable in their contexts to aid their work. Last, but not least, we hope that researchers in the areas of economics of innovation and science of science will find here sources for fresh ideas in these important areas.
@@
Economic studies that attempt to draw a general link between research and economic benefit go back more than a century, to the 1912 publication of Schumpeter’s Theorie der wirtschaftlichen Entwicklung (Theory of Economic Development) [17]. Qualitative analyses have long pointed out relationships between investment in research and economic growth. A current example of this sort of approach, used in the USA since at least the 1990s, can be found in [18]. This sort of analysis links broad areas of economic activity, such as telecommunications, back to specific suites of USA federal government investment in research. While these qualitative relationships are convincing, they are drawn out only at very large scales and only in rare cases point back to a single research-enabling facility as a causal factor. As a result, such approaches are not generally helpful in evaluating an individual research-enabling facility and not at all helpful in managing budgets within institutions of higher education.
@@
In 2016, Del Bo reviewed the rate of return of large-scale “research infrastructures,” which is a term from the literature in the European Union [20]. The term “research infrastructure” is derived from Florio’s early work and is defined as “high-capital intensity and long-lasting facilities and equipment, typically operating in oligopoly conditions, whose objective is to support economic development and produce social benefits through the generation of new knowledge and, often, other spillover effects” [21]. Del Bo’s analysis of return on investment (ROI) in research infrastructures helpfully creates a categorization of studies based on scope, ranging from macroeconomic to project-level studies [20]. Del Bo’s excellent review of macroeconomic methods applicable to research infrastructures is based largely on endogenous growth models. Del Bo’s efforts were focused on the European Union, but the methods outlined are broadly applicable to any situation involving an appropriately large and long-lasting research infrastructure.
The European facilities of CERN (Conseil européen pour la recherche nucléaire - the Europe-based organization for Nuclear Research) - is perhaps the best-known example of a persistent and large-scale Research Infrastructure in the sense of Del Bo’s definition. The investments in CERN are so large that it is possible to assess their impacts through national and multinational macroeconomic studies. Florio followed his early paper-length analyses with an extensive book devoted to “social cost-benefit of research infrastructures” [22]. This book offers a thorough conceptual overview of types of benefit that may derive from investment in large-scale Research Infrastructures and quantitative methods for estimating such effects. In summary, research into the public value or public investment in CERN shows it has a net positive impact on society. Many other of the Research Infrastructures reviewed by Del Bo also have a net positive economic impact.
@@
CERN and cyberinfrastructure facilities at universities in the USA have been the subject of more cost-benefit and ROI analyses than most - perhaps any - research-enabling facilities in the world. Most cyberinfrastructure facilities involve orders of magnitude less investment and less economic impact. As a result, their impact is generally not detectable via the macroeconomic approaches used in analysis of CERN and other large-scale Research Infrastructures. Recognizing the extensive use of the term Research Infrastructure as defined by Florio and implying persistent and very large-scale facilities, we will focus on activities at a smaller scale. We use the term research-enabling facility, as the name implies, to signify facilities at scales smaller than a Research Infrastructure as defined by Florio.
In this section, we present project-level approaches to quantitative assessment of investment in cyberinfrastructure. Much of the research cited in this current paper is project-level analysis of one of the largest cyberinfrastructure projects ever funded by the USA’s National Science Foundation (NSF), called XSEDE - the eXtreme Science and Engineering Discovery Environment. XSEDE included supercomputers, visualization resources, data storage facilities, and training and support services, described in detail in [26] and [11]. XSEDE operated from 2011 to 2022, providing services supporting non-classified research by the USA national research community. Because of its scale and visibility, the value provided to the USA by XSEDE has been examined with notable care.
Our discussion of project-level approaches to evaluating research-enabling facilities uses a categorization based on that presented by Del Bo and expanded to include other and newer analytic methods:
@@
Note that in every area save two, our discussion includes research results by one or more of the authors on this paper except: discussion of time savings, which we think is interesting and bears mention even though we have not done this sort of work ourselves; and bibliometric analyses, an area in which work has been done under subcontract from coauthor Towns.
We focus on the critical information that potentially skeptical stakeholders, including government officials and the general public, commonly ask for: Is the work enabled by a research-enabling facility sufficiently financially significant that the investment in that facility is financially worthwhile and beneficial to society as a whole?
@@
The total quality management movement reached the USA with force in the 1980s. Total quality management is “a term first used to describe a management approach to quality improvement” [27] and is explained in detail in [28]. One focus of total quality management is the creation and measurement of a set of metrics that meaningfully define quality and costs for an organization. Such metrics have long included client or user reported satisfaction, total cost of ownership (TCO), and usage metrics that are typically specific to organizational type and an organization’s specific goals. Such approaches have been applied to computing services in academia since at least the 1990s [29]. All of these metrics are useful, but they have limitations.
Client satisfaction surveys are a basic aspect of the quality assessment of any research-enabling infrastructure or facility. (Term usage varies. “Client” is regarded as a respectful term among many who are steeped in TQM methodologies, although the term “user” is also commonly used.) Of course, highly negative survey ratings indicate problems. Positive ratings demonstrate that the users of a facility like it, which is good. However, positive satisfaction scores do not necessarily prove that a facility is a cost-effective or strategic investment [29].
@@
For many categories of research-enabling facility there are also specific sorts of traditional operational metrics. A specific organization may also have certain goals tied to strategic objectives. Typical metrics for the operation of cyberinfrastructure facilities include percentage uptime (time during which a system was available for use as a percentage of theoretical maximum), the number of jobs submitted and completed, CPU- and GPU- hours used per unit, and the amount of data stored. Again, all of these metrics must be carefully interpreted. Unusually low usage numbers relative to resource availability may be indicative of some sort of problem - perhaps there has been insufficient training, or the system is not perceived by users to match their needs. Low usage statistics may also be found immediately after a large upgrade to the capability of computational systems when the people who use the system have not adjusted to the increased availability of cycles. Consistent usage near peak capacity may be indicative of a balanced match of availability and need. Consistent usage at peak capacity may indicate inadequate supply, or a doctoral student iteratively reanalyzing their data to avoid actually writing their dissertation, or a faculty member running mindless parameter sweeps rather than running carefully thought out and efficient computational experiments. (These are not hypothetical examples; the last named author was responsible for HPC systems for 30 years and experienced every one of these examples at least once.)
There will be discipline-specific operational metrics appropriate to any sort of research-enabling facility. While such metrics can clarify the activities and utilization of a research-enabling facility among practitioners in a given area, these operational metrics are not easily translatable into terms that relate value to investment. As such, they are unlikely to resonate intuitively with certain important categories of stakeholders, potentially ranging from taxpaying voters to elected officials to University Presidents and CFOs (Chief Financial Officers).
@@
One example of analysis centered on research facilities is presented in von Laszewski et al. [32], who performed a bibliometric analysis of peer-reviewed scientific publications created with analyses of simulations performed on XSEDE cyberinfrastructure facilities. They investigated the question: Are research publications created using XSEDE cited more often than similar papers created without the use of XSEDE? Their analysis suggested that the answer was yes. Unfortunately, however, this analysis suffered from a significant conceptual shortcoming. To investigate the impact of a specific research-enabling facility on overall research output, one really ought to ask two questions: 1) Are papers produced using a certain type of research-enabling facility more impactful, in general, than similar papers produced without this type of research-enabling facility? 2) Among papers produced using one type of research-enabling facility, are those produced using one specific facility (or perhaps one set of instruments) more impactful than other papers produced using facilities of the same type (or instruments of the same category)? The line of analysis suggested by von Laszewski et al. is interesting, but a more careful analysis is needed, hopefully published in one of the standard journals that focus on bibliometrics.
@@
Empirically derived econometric estimators can be used to estimate the direct and indirect economic effects of investments in many kinds of projects and services. The Regional Input–Output Modeling System (RIMSII) uses empirically derived multipliers to estimate the total direct and indirect economic impact and the total number of jobs created by a variety of investments [34]. These multipliers are calculated for a variety of investments, ranging from building a parking garage to investing in jobs and facilities at a university, for specific geographic regions. IMPLAN (originally short for “impact analysis for planning”) offers the same approach globally [35]. Because these approaches use general categories of investment, they may underestimate the value of investment in highly unusual areas. For example, when estimating the value of investment in advanced cyberinfrastructure systems within the USA, one could potentially use a RIMSII multiplier from one of two types of “Professional, scientific, and technical services,” or one of two types of “Junior colleges, colleges, universities, and professional schools.” The impact of some specialized research-enabling facility might well be greater than the calculated average for such relatively broad categories. However, these estimation methods are well known and widely trusted. They produce figures that constitute an unimpeachable lower bound for the overall economic impact of any investment. This approach has been used to demonstrate positive impact of a cyberinfrastructure center on its home state [36].
@@
One of the earliest concepts in accounting for the value of a particular instance of publicly funded research-enabling facility has been referred to as “leverage.” For publicly funded research facilities, leverage is simply the financial value of the sum of grant awards received by the publicly funded research projects that make use of that facility. This total value can be useful in concise statements that explain the value of a facility. Using data from [8] about how much the USA federal government invested in XSEDE and how much it invested in other grant awards that used XSEDE resources, for example, one could say that “The NSF leveraged USD 200M of investment in XSEDE to support USD 4.5B in total grant awards that made use of XSEDE facilities and services.” Indeed, this sort of statement has often been used by coauthor Towns in non-published materials prepared for the NSF as part of XSEDE project reviews. However, these figures must be used with a measure of care because of the matter of shared credit. Suppose that a discovery critically depends on two different research-enabling facilities. The value of such a discovery can be considered as “leverage” for each of those facilities. But it would be double counting to count the full value as part of the “return on the investment” in each facility, something that has been done in some reports. It is certainly unlikely in general to be wholly true that research conducted with the use of a particular facility could or would not have been successfully carried out without the use of that facility. Such a claim might be true in the case of the research conducted with the Large Hadron Collider that led to the verification of the existence of the Higgs Boson. But, in general, one is well advised to be wary of contrapositive statements of the form “without research-enabling facility Z, discovery Beta would not have been possible.” Certainly in the case of many kinds of research-enabling facilities, some of the research done using that facility might well have been successfully carried out by clever scientists using some other facility. Some other research might have occurred, but not as efficiently or as well. And yet other research may indeed not have happened at all. So, while a “leverage” total can be useful shorthand, it is necessary to be careful when using this term.
@@
This formula is essentially a way to calculate the financial efficiency of providing a service. In terms of public sector research and university activities, this formula shows whether something is being done at greater or lesser cost than would be experienced if products or services were simply bought on the open market. ROIproxy is a way to measure the financial effectiveness resulting from investments in cyberinfrastructure over any time period, in the absence of anything being bought or sold. In this formula, as in the formula for ROI, a value of 1.0 indicates a break-even level. A value greater than 1.0 indicates that the value of services provided when calculated based on market costs exceeds the actual cost of operating the facility. These calculations are conceptually straightforward. The denominator is the TCO for the relevant period of time. For the numerator, one tallies as many services and products as possible that were created using the cyberinfrastructure resources, finds the prevailing market prices for as many of these services as possible, and arrives at a total estimated market value for the relevant time period. Credible valuations are available for a surprising array of entities, including researcher time, scientific accomplishments such as a peer-reviewed publication or Nobel Prize, educational outcomes, and even improvements in quality of life. These valuations may be done annually or over longer time periods. The results are approximate, of course. Calculated results are likely to be an underestimation of the actual long-term value. There will often be benefits that are impossible to quantify and thus are not included in the calculations. Further, the economic value of some benefits will increase over time as the scope of application expands. Examples include health care and other benefits to people’s quality of life.
@@
A value >1 favors on-premises as the more economical choice. Note that cost-benefit analyses are very similar and in general each of these two forms of analysis may be straightforwardly restated in the terminology of the other form of analysis.
@@
The International Integrated Reporting Framework is of use in answering this more difficult and far more important question. The purpose of the International $<$IR$>$ Framework (as it is typically referred to) is “to explain to providers of financial capital how an organization creates, preserves, or erodes value over time” [38]. This framework includes six categories, intended to capture all possible types of value creation: financial capital, manufactured capital, intellectual capital, human capital, social and relationship capital, and natural capital. This structure, according to the IFRS (International Financial Reporting Standards) website, benefits all stakeholders, including “legislators, regulators, and policymakers.” The IFRS has a grand goal of transforming corporate reporting, so the description of their reporting framework extends well beyond what is needed for an assessment of the value of an instance of a research-enabling facility. However, the comprehensiveness of the framework, its clarity, and its backing and promotion by an international consortium of accountancy researchers and experts provide great credibility to the portion relevant to the evaluation of research facilities and research projects generally (pp. 18-23 in [38]).
@@
Within academia, it can be useful to quantitatively examine important outputs even if it is not immediately possible to associate a financial value with them. And there are indeed, non-financial metrics within academia that are important in an institution’s reputation, ranking, and its recruitment of students, faculty, and staff. In a 1928 work, Cobb and Douglas proposed a “production function” that expressed factory output (Y) as a function of investment in capital (K) and labor (L) [39]. Smith built on this approach to relate the production of a variety of academic outputs (Y) to capital investment in systems (as measured in TeraFLOPS) and labor costs for the staff that operate and support them as[40]:
@@
Table 1: Relationship of academic output metrics to on-premises HPC TeraFLOPS capacity and investment in staff supporting use of HPC systems at Purdue University, 1999–2020.
@@
Table 2: Percent variance in academic output metrics at Purdue University, explained by on-premises HPC TeraFLOPS capacity and investment in staff supporting use of HPC systems at Purdue University, 1999–2020.
@@
Table 3: Relationship of academic output metrics to on-premises HPC TeraFLOPS capacity and investment in staff supporting use of HPC systems at 5 R1 universities.
@@
Table 4: Percent variance in academic output metrics explained by on-premises HPC TeraFLOPS capacity and investment in staff supporting use of HPC systems at 5 R1 universities.